# AVOCADO: Adaptive Optimal Collision Avoidance driven by Opinion

### v1

@@ IV-B Nonlinear Opinion Dynamics Adaptive Law
$x_{i}=2\alpha_{i}-1\Leftrightarrow\alpha_{i}=\frac{x_{i}+1}{2}.$ (9)
The purpose of this change is to shift the degree of cooperation from the interval $[0,1]$ to the interval $[-1,1]$, which better suits the definition of nonlinear opinion dynamics provided in Eq. (5). From now on, we refer to $x_{i}$ as the shifted degree of cooperation of agent $i$ with respect to the robot.
@@
$\dot{x}_{i}=-d_{i}x_{i}+d_{i}A_{i}\mathsf{tanh}\left(a_{i}x_{i}+c_{i}y_{i}\right)+b_{i}.$ (10)
Compared to (5), Eq. (10) simplifies the consensus term to a weighted sum of the current shifted degree of cooperation $x_{i}$ and a quantity $y_{i}$ that will be defined later and which represents the opinion would the agent $i$ have about $x_{i}$. Ideally, the consensus term would include not only $y_{i}$, but also all the other agents. However, this would require some sort of communication infrastructure or $k$-hop neighboring dependence that is not feasible in a real setting that only depends on onboard sensing. Therefore, we assume that the opinion of the robot on the degree of cooperation of agent $i$ is a pair-wise interaction isolated from other agents. In this sense, the design of (10) manifests a fundamental difference with respect to existing nonlinear opinion dynamics formulations, tailored towards a feasible implementation in a real robotic platform. The quantities $a_{i},b_{i},c_{i},d_{i}$ are gains with similar meaning to those in Eq. (5) and which parameterized the adaptive law. For instance, $b_{i}=-0.5$ represents some inductive bias on the degree of cooperation of agent $i$ towards not cooperating with the robot. As nonlinear function, we choose $\mathsf{tanh}$ to respect the domain of $x_{i}$. Finally, $A_{i}$ represents the attention, and is a variable that dynamically evolves with time, in contrast to the algebraic relationship $g(x_{j})$ formulated in Eq. (5). As depicted in Fig. 3, the idea is to design and tune (10) such that, when the agent $i$ is close to the robot and does not collaborate to avoid the collision, the (shifted) degree of cooperation associated to agent $i$ evolves such that the robot takes the responsibility. On the other hand, if agent $i$ modifies its velocity to avoid collision, then the (shifted) degree of cooperation evolves to enable the robot a more selfish behavior. The attention term $A_{i}$ appears scaled in (10) by gain $d_{i}$. The reason behind this scaling will be explained later, in Section IV-C.
@@
There are two variables in Eq. (10) that are not readily available at the robot. The first one is the attention, which should capture the importance of adapting the shifted degree of cooperation or if, otherwise, the shifted degree of cooperation must not change. The second quantity is $y_{i}$, which a priori depends on the information at agent $i$. We will later prove how the robot can estimate $y_{i}$ from just onboard sensor observations. Now, we focus on the attention $A_{i}$.
@@
The next step is to define $y_{i}$ and develop a method to compute it from sensor observations. According to the definition of the nonlinear opinion dynamics in Eq. (10), $y_{i}$ represents the opinion that agent $i$ has on the value of the shifted degree of cooperation $x_{i}$. Recalling the reasoning behind the nonlinear opinion dynamics design in (5), the consensus term $a_{i}x_{i}+c_{i}y_{i}$ aims at fusing the values that the robot and agent $i$ have on the degree of cooperation $x_{i}$. In the case of the robot, this is simply the current value of $x_{i}$; in the case of agent $i$, this is precisely the the opinion that agent $i$ has on the value of the shifted degree of cooperation $x_{i}$. In typical opinion dynamics formulations it is assumed that these values can be communicated. This is not the case in this work, so we need to recover $y_{i}$ using an estimation method that relies solely on onboard sensing.
@@
Figure 4: Geometry behind the projection estimator for ${y}_{i}$.
@@
These insights motivate the following projection estimator of $y_{i}$:
${y}_{i}=\mathsf{tanh}\left(\varepsilon\left(\frac{||\mathsf{proj}(\Delta\mathbf{v}_{i},\mathbf{u}_{i})||}{||\mathbf{u}_{i}||}-\frac{1}{2}\right)\right).$ (15)
The projection operator $\mathsf{proj}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\mathbf{b}\cdot\mathbf{b}}\mathbf{b}$ projects the vector $\Delta\mathbf{v}_{i}$ on vector $\mathbf{u}_{i}$. Furthermore, by normalizing with the norm $||\mathbf{u}_{i}||$, the estimate of $\alpha_{i}$, $\hat{\alpha}_{i}$, is extracted. The other operations translate the estimated degree of cooperation to the estimated shifted degree of cooperation ${y}_{i}$, where $\varepsilon>0$ tunes how sensitive is the $\mathsf{tanh}$ function to changes in $\Delta\mathbf{v}_{i}$. In this sense, we remark that Eq. (15) is not a perfect estimator since it only relies on the sensor observations of the change in velocities $\Delta\mathbf{v}_{i}$. This motivates the use of the function $\mathsf{tanh}$, which simultaneously restricts the domain of the estimation to $[-1,1]$ and allows a fast evolution of the estimated shifted degree of cooperation. For instance, if the robot observes that agent $i$ moves straight to it and $\Delta\mathbf{v}_{i}\approx 0$, then this means that ${y}_{i}\approx-1$ and the robot estimates that agent $i$ is not cooperating.
@@ IV-C Nonlinear Opinion Dynamics Adaptive Law Tuning
First, we examine the stability properties of the adaptive law and the conditions under which bifurcation, i.e., instability of the initially stable equilibrium point of Eq. (10), happens. To do so, we use the expression $\text{sech}(x)=1/\cosh(x)$.
@@
Let $d_{i}>0$, $a_{i}>0$ and $b_{i}\in[-1,1]$. Then, $x_{i}=b_{i}/d_{i}$ is an unstable equilibrium point of the nonlinear opinion dynamics adaptive law in (10) if $A_{i}>1/\left(a_{i}\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)\right)$.
@@
Initially, $A_{i}=0$, ${y}_{i}=0$ and $\dot{x}_{i}=0$. Therefore, the robot has an initial opinion $x_{i}=b_{i}/d_{i}$ that is an equilibrium of the nonlinear opinion dynamics adaptive law in Eq. (10). Nevertheless, whether this equilibrium is stable or unstable depends on the parameters of the adaptive law. Let
@@
be the unique eigenvalue of (10) after linearization at $x_{i}=b_{i}/d_{i}$. Equilibrium becomes unstable when $\lambda>0$. Assuming that $a_{i}$, $d_{i}$ and $b_{i}$ are fixed, and taking into account that $\text{sech}(x)\in(0,1]$ in all its domain, $x_{i}=b_{i}/d_{i}$ becomes an unstable equilibrium when $A_{i}>1/\left(a_{i}\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)\right)$. ∎
@@
If $b_{i}=0$, then $\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)=1$ and the condition on the attention level reduces to $A_{i}>1/a_{i}$. Since $\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)\in(0,1)$, $A_{i}>1/a_{i}$ is an upper bound on the attention value beyond which the equilibrium $x_{i}=b_{i}/d_{i}$ becomes an unstable equilibrium and, therefore, $A_{i}>1/a_{i}$ is a sufficient condition for $x_{i}=b_{i}/d_{i}$ being an unstable equilibrium.
@@
$x_{i}^{\mathsf{eq}}=A_{i}\mathsf{tanh}(a_{i}x_{i}^{\mathsf{eq}}+c_{i}{y}_{i})+\frac{b_{i}}{d_{i}}.$ (17)
Besides, $|x_{i}^{\mathsf{eq}}|\in\left[\frac{|b_{i}|}{d_{i}}-1,\frac{|b_{i}|}{d_{i}}+1\right]$.
@@
From the assumption that $A_{i}>\frac{1}{a_{i}}$ holds, in the transition $A_{i}=\frac{1}{a_{i}}$ bifurcation happens [69]. The dynamics in Eq, (10) are bounded due to the term $-d_{i}x_{i}$, so equilibrium(s) point(s) exist(s) after bifurcation. By enforcing the equilibrium condition $\dot{x}_{i}=0$ in Eq. (10), Eq. (17) is obtained. We recall that $A_{i}\in[0,1]$ and $|\mathsf{tanh}(\bullet)|\in[0,1]$. Therefore, the absolute value of the equilibrium $|x_{i}^{\mathsf{eq}}|$ is bounded and such that $|x_{i}^{\mathsf{eq}}|\in\left[\frac{|b_{i}|}{d_{i}}-1,\frac{|b_{i}|}{d_{i}}+1\right]$. ∎
@@
If there is no bias, $b_{i}=0$ and $x_{i}^{\mathsf{eq}}\in\left[-1,1\right]$; otherwise, the bias introduces a shift in the domain that naturally encodes a priori knowledge on the degree of cooperation of the agent. Proposition 2 motivates the use of the scaling factor $d_{i}$ in the attention $A_{i}$: in the absence of this factor, the expression of the equilibrium points is $x_{i}^{\mathsf{eq}}=\frac{A_{i}\mathsf{tanh}(a_{i}x_{i}^{\mathsf{eq}}+c_{i}{y}_{i})+b_{i}}{d_{i}}$, leading to an asymmetrical domain for $|x_{i}^{\mathsf{eq}}|$ with respect to $b_{i}/d_{i}$, i.e., not symmetrical with respect to the initial equilibrium before bifurcation. This asymmetry would derive in an intrinsic bias towards any of the two subsequent equilibrium points, which is an undesired effect that we overcome by including the scaling factor $d_{i}$.
@@
Finally, $c_{i}$ shall be designed to guarantee that the nonlinear term $\mathsf{tanh}(\bullet)$ in the adaptive law is enough sensitive to the behavior of agent $i$, reacting fast enough to avoid collisions. On the other hand, $\varepsilon$ in (15) tunes how sensitive is the estimator of ${y}_{i}$ to changes in the speed of agent $i$.
@@
In summary, the nonlinear opinion dynamics adaptive law depends on six parameters: (i) $\varepsilon>0$ determines how sensitive is the estimator of ${y}_{i}$ and should be chosen large enough to ensure ${y}_{i}\approx 1$ with enough time to avoid collision but small enough to avoid noisy estimations; (ii) $\kappa_{i}>0$ and $\delta_{i}\in[0,1)$ determines how fast the adaptive law approaches the bifurcation point, and therefore establishes when the robot decides its opinion about the degree of cooperation of agent $i$; (iii) $d_{i}>0$ sets the convergence speed of the dynamics of the adaptive law, associated to the forgetting factor $-d_{i}x_{i}$; (iv) $a_{i}$ determines the attention level necessary to reach the bifurcation, taking into account the value of $d_{i}$; (v) $b_{i}$ is a bias that encodes potential prior knowledge on the (shifted) degree of cooperation of agent $i$; (vi) $c_{i}$ weights the importance of the estimate ${y}_{i}$ compared to the current value of $a_{i}x_{i}$ such that, if $a_{i}>>c_{i}$, the robot tends to preserve its opinion over the (shifted) degree of cooperation, or heavily relies on the estimator if $c_{i}>>a_{i}$.
IV-D Exploiting attention to overcome deadlocks

### v4

@@ IV-B Nonlinear Opinion Dynamics Adaptive Law
${o}_{i}=2\alpha_{i}-1\Leftrightarrow\alpha_{i}=\frac{{o}_{i}+1}{2}.$ (9)
The purpose of this change is to shift the degree of cooperation to better suit the definition of nonlinear opinion dynamics provided in Eq. (5). From now on, we refer to ${o}_{i}$ as the shifted degree of cooperation of agent $i$ with respect to the robot.
@@
$\dot{{o}}_{i}=-d_{i}{o}_{i}+d_{i}A_{i}\mathsf{tanh}\left(a_{i}{o}_{i}+c_{i}{e}_{i}\right)+b_{i}.$ (10)
Compared to (5), Eq. (10) simplifies the consensus term to a weighted sum of the current shifted degree of cooperation ${o}_{i}$ and a quantity ${e}_{i}$ that will be defined later and which represents an estimate of the opinion agent $i$ has about ${o}_{i}$. Ideally, the consensus term would include not only ${e}_{i}$, but also all the other agents. However, this would require some sort of communication infrastructure or $k$-hop neighboring dependence that is not feasible in a real setting that only depends on onboard sensing. Therefore, we assume that the opinion of the robot on the degree of cooperation of agent $i$ is a pair-wise interaction isolated from other agents. In this sense, the design of (10) manifests a fundamental difference with respect to existing nonlinear opinion dynamics formulations, tailored towards a feasible implementation in a real robotic platform. The quantities $a_{i},b_{i},c_{i},d_{i}$ are gains with similar meaning to those in Eq. (5) and which parameterized the adaptive law. For instance, $b_{i}=-0.5$ represents some inductive bias on the degree of cooperation of agent $i$ towards not cooperating with the robot. As nonlinear function, we choose $\mathsf{tanh}$ to respect the domain of ${o}_{i}$. Finally, $A_{i}$ represents the attention, and is a variable that dynamically evolves with time, in contrast to the algebraic relationship $g({o}_{j})$ formulated in Eq. (5). As depicted in Fig. 4, the idea is to design and tune (10) such that, when the agent $i$ is close to the robot and does not collaborate to avoid the collision, the (shifted) degree of cooperation associated to agent $i$ evolves such that the robot takes the responsibility. On the other hand, if agent $i$ modifies its velocity to avoid collision, then the (shifted) degree of cooperation evolves to enable the robot a more selfish behavior. The attention term $A_{i}$ appears scaled in (10) by gain $d_{i}$. The reason behind this scaling will be explained later, in Section IV-C.
@@
There are two variables in Eq. (10) that are not readily available at the robot. The first one is the attention, which should capture the importance of adapting the shifted degree of cooperation or if, otherwise, the shifted degree of cooperation must not change. The second quantity is ${e}_{i}$, which a priori depends on the information at agent $i$. We will later prove how the robot can estimate ${e}_{i}$ from just onboard sensor observations. Now, we focus on the attention $A_{i}$.
@@
The next step is to define ${e}_{i}$ and develop a method to compute it from sensor observations. According to the definition of the nonlinear opinion dynamics in Eq. (10), ${e}_{i}$ represents the opinion that agent $i$ has on the value of the shifted degree of cooperation ${o}_{i}$. Recalling the reasoning behind the nonlinear opinion dynamics design in (5), the consensus term $a_{i}{o}_{i}+c_{i}{e}_{i}$ aims at fusing the values that the robot and agent $i$ have on the degree of cooperation ${o}_{i}$. In the case of the robot, this is simply the current value of ${o}_{i}$; in the case of agent $i$, this is precisely the opinion that agent $i$ has on the value of the shifted degree of cooperation ${o}_{i}$. In typical opinion dynamics formulations it is assumed that these values can be communicated. This is not the case in this work, so we need to recover ${e}_{i}$ using an estimation method that relies solely on onboard sensing.
@@
Figure 5: Geometry behind the projection estimator for ${e}_{i}$.
@@
These insights motivate the following projection estimator of ${e}_{i}$:
${e}_{i}=\mathsf{tanh}\left(\varepsilon\left(\frac{||\mathsf{proj}(\Delta\mathbf{v}_{i},\mathbf{u}_{i})||}{||\mathbf{u}_{i}||}-\frac{1}{2}\right)\right).$ (15)
The projection operator $\mathsf{proj}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\mathbf{b}\cdot\mathbf{b}}\mathbf{b}$ projects the vector $\Delta\mathbf{v}_{i}$ on vector $\mathbf{u}_{i}$. Furthermore, by normalizing with the norm $||\mathbf{u}_{i}||$, the estimate of $\alpha_{i}$ is extracted. The other operations translate the estimated degree of cooperation to the estimated shifted degree of cooperation ${e}_{i}$, where $\varepsilon>0$ tunes how sensitive is the $\mathsf{tanh}$ function to changes in $\Delta\mathbf{v}_{i}$. In this sense, we remark that Eq. (15) is not a perfect estimator since it only relies on the sensor observations of the change in velocities $\Delta\mathbf{v}_{i}$. This motivates the use of the function $\mathsf{tanh}$, which simultaneously restricts the domain of the estimation to $[-1,1]$ and allows a fast evolution of the estimated shifted degree of cooperation. For instance, if the robot observes that agent $i$ moves straight to it and $\Delta\mathbf{v}_{i}\approx 0$, then this means that ${e}_{i}\approx-1$ and the robot estimates that agent $i$ is not cooperating. It is worth noting that the projection estimator in (15) directly uses the velocity measurements from onboard sensors, which are subject to disturbances. However, designing the adaptive law as a non-linear opinion dynamic naturally provides robustness against small perturbations, and exploits past information in the opinion estimation by means of the memory term $-d_{i}{o}_{i}$ [73]. Therefore, the proposed adaptive law can be framed as a nonlinear estimator with a dynamic model (Eq. (10)) and a measurement model (Eq. (15)). The sensitivity and robustness of the general formulation of the nonlinear opinion dynamics model is analyzed in previous works [72]. The next section analyzes specific considerations on how to tune the parameters of AVOCADO.
@@ IV-C Nonlinear Opinion Dynamics Adaptive Law Tuning
First, we examine the stability properties of the adaptive law and the conditions under which bifurcation, i.e., instability of the initially stable equilibrium point of Eq. (10), happens. To do so, we use the expression $\text{sech}(\bullet)=1/\cosh(\bullet)$.
@@
Let $d_{i}>0$, $a_{i}>0$ and $b_{i}\in[-1,1]$. Then, ${o}_{i}=b_{i}/d_{i}$ is an unstable equilibrium point of the nonlinear opinion dynamics adaptive law in (10) if $A_{i}>1/\left(a_{i}\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)\right)$.
@@
Initially, $A_{i}=0$, ${e}_{i}=0$ and $\dot{{o}}_{i}=0$. Therefore, the robot has an initial opinion ${o}_{i}=b_{i}/d_{i}$ that is an equilibrium of the nonlinear opinion dynamics adaptive law in Eq. (10). Nevertheless, whether this equilibrium is stable or unstable depends on the parameters of the adaptive law. Let
@@
be the unique eigenvalue of (10) after linearization at ${o}_{i}=b_{i}/d_{i}$. Equilibrium becomes unstable when $\lambda>0$. Assuming that $a_{i}$, $d_{i}$ and $b_{i}$ are fixed, and taking into account that $\text{sech}(\bullet)\in(0,1]$ in all its domain, ${o}_{i}=b_{i}/d_{i}$ becomes an unstable equilibrium when $A_{i}>1/\left(a_{i}\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)\right)$. ∎
@@
If $b_{i}=0$, then $\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)=1$ and the condition on the attention level reduces to $A_{i}>1/a_{i}$. Since $\text{sech}^{2}\left(a_{i}\frac{b_{i}}{d_{i}}\right)\in(0,1)$, $A_{i}>1/a_{i}$ is an upper bound on the attention value beyond which the equilibrium ${o}_{i}=b_{i}/d_{i}$ becomes an unstable equilibrium and, therefore, $A_{i}>1/a_{i}$ is a sufficient condition for ${o}_{i}=b_{i}/d_{i}$ being an unstable equilibrium.
@@
${o}_{i}^{\mathsf{eq}}=A_{i}\mathsf{tanh}(a_{i}{o}_{i}^{\mathsf{eq}}+c_{i}{e}_{i})+\frac{b_{i}}{d_{i}}.$ (17)
Besides, $|{o}_{i}^{\mathsf{eq}}|\in\left[\frac{|b_{i}|}{d_{i}}-1,\frac{|b_{i}|}{d_{i}}+1\right]$.
@@
From the assumption that $A_{i}>\frac{1}{a_{i}}$ holds, in the transition $A_{i}=\frac{1}{a_{i}}$ bifurcation happens [72]. The dynamics in Eq, (10) are bounded due to the term $-d_{i}{o}_{i}$, so equilibrium(s) point(s) exist(s) after bifurcation. By enforcing the equilibrium condition $\dot{{o}}_{i}=0$ in Eq. (10), Eq. (17) is obtained. We recall that $A_{i}\in[0,1]$ and $|\mathsf{tanh}(\bullet)|\in[0,1]$. Therefore, the absolute value of the equilibrium $|{o}_{i}^{\mathsf{eq}}|$ is bounded and such that $|{o}_{i}^{\mathsf{eq}}|\in\left[\frac{|b_{i}|}{d_{i}}-1,\frac{|b_{i}|}{d_{i}}+1\right]$. ∎
@@
If there is no bias, $b_{i}=0$ and ${o}_{i}^{\mathsf{eq}}\in\left[-1,1\right]$; otherwise, the bias introduces a shift in the domain that naturally encodes a priori knowledge on the degree of cooperation of the agent. Proposition 2 motivates the use of the scaling factor $d_{i}$ in the attention $A_{i}$: in the absence of this factor, the expression of the equilibrium points is ${o}_{i}^{\mathsf{eq}}=\frac{A_{i}\mathsf{tanh}(a_{i}{o}_{i}^{\mathsf{eq}}+c_{i}{e}_{i})+b_{i}}{d_{i}}$, leading to an asymmetrical domain for $|{o}_{i}^{\mathsf{eq}}|$ with respect to $b_{i}/d_{i}$, i.e., not symmetrical with respect to the initial equilibrium before bifurcation. This asymmetry would derive in an intrinsic bias towards any of the two subsequent equilibrium points, which is an undesired effect that we overcome by including the scaling factor $d_{i}$.
@@
Finally, $c_{i}$ shall be designed to guarantee that the nonlinear term $\mathsf{tanh}(\bullet)$ in the adaptive law is enough sensitive to the behavior of agent $i$, reacting fast enough to avoid collisions. On the other hand, $\varepsilon$ in (15) tunes how sensitive is the estimator of ${e}_{i}$ to changes in the speed of agent $i$.
@@
In summary, the nonlinear opinion dynamics adaptive law depends on six parameters: (i) $\varepsilon>0$ determines how sensitive is the estimator of ${e}_{i}$ and should be chosen large enough to ensure ${e}_{i}\approx 1$ with enough time to avoid collision but small enough to avoid noisy estimations; (ii) $\kappa_{i}>0$ and $\delta_{i}\in[0,1)$ determines how fast the adaptive law approaches the bifurcation point, and therefore establishes when the robot decides its opinion about the degree of cooperation of agent $i$; (iii) $d_{i}>0$ sets the convergence speed of the dynamics of the adaptive law, associated to the forgetting factor $-d_{i}{o}_{i}$; (iv) $a_{i}$ determines the attention level necessary to reach the bifurcation, taking into account the value of $d_{i}$; (v) $b_{i}$ is a bias that encodes potential prior knowledge on the (shifted) degree of cooperation of agent $i$; (vi) $c_{i}$ weights the importance of the estimate ${e}_{i}$ compared to the current value of $a_{i}{o}_{i}$ such that, if $a_{i}>>c_{i}$, the robot tends to preserve its opinion over the (shifted) degree of cooperation, or heavily relies on the estimator if $c_{i}>>a_{i}$.
IV-D Exploiting attention to avoid symmetry deadlocks